OpenStudy (deadshot):

How do I write 5+7+9+11+13+15+17 in sigma notation?

OpenStudy (anonymous):

What is the common factor that we are adding each time?

OpenStudy (deadshot):

2

OpenStudy (anonymous):

Right, so we're looking at something that looks like: \[\sum Number + 2n\] The "Number" will depend on where you want to start your summation. Usually, we begin from n = 0 or from n = 1. We'll need to make that choice now. Which would you like to do?

OpenStudy (deadshot):

n=0

OpenStudy (anonymous):

Alright, so if n = 0, and we want to get the number 5, what is our "Number?"

OpenStudy (deadshot):

would it be 3?

OpenStudy (anonymous):

Not 3 We want: Number + 2*0 = 5

OpenStudy (deadshot):

so, it would be 5? since 2*0=0

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

So, we now we just need to decide when to stop our summation. The last term that we want is 17. So, what is n when the result is 17? 5+2n = 17

OpenStudy (deadshot):

6?

OpenStudy (anonymous):

Yep :) And we just put it all together: \[\sum_{n=0}^{6}5+2n = 5+7+9+11+13+15+17\]

OpenStudy (deadshot):

Thanks!

OpenStudy (anonymous):

Does it all make sense?

OpenStudy (deadshot):

yes

OpenStudy (anonymous):

Outstanding :)